# WikiReddit: Tracing Information and Attention Flows Between Online Platforms

### v1

@@ Abstract
The World Wide Web is a complex interconnected digital ecosystem, where information and attention flow between platforms and communities throughout the globe. These interactions co-construct how we understand the world, reflecting and shaping public discourse. Unfortunately, researchers often struggle to understand how information circulates and evolves across the web because platform-specific data is often siloed and restricted by linguistic barriers. To address this gap, we present a comprehensive, multilingual dataset capturing all Wikipedia links shared in posts and comments on Reddit from 2020 to 2023, excluding those from private and NSFW subreddits. Each linked Wikipedia article is enriched with revision history, page view data, article ID, redirects, and Wikidata identifiers. Through a research agreement with Reddit, our dataset ensures user privacy while providing a query and ID mechanism that integrates with the Reddit and Wikipedia APIs. This enables extended analyses for researchers studying how information flows across platforms. For example, Reddit discussions use Wikipedia for deliberation and fact-checking which subsequently influences Wikipedia content, by driving traffic to articles or inspiring edits. By analyzing the relationship between information shared and discussed on these platforms, our dataset provides a foundation for examining the interplay between social media discourse and collaborative knowledge consumption and production.
@@ Introduction
In addition to Wikipedia, audiences consult a range of online services for news including news websites, apps, and social media (St. Aubin and Liedke 2024; Vraga and Tully 2021). People who include, but do not exclusively rely on, social media in their news diets tend to have higher news media knowledge (Schulz, Fletcher, and Nielsen 2024). Among these social media sites is Reddit, a social media platform where community members share, vote, and comment on content and foster community-based engagement on topics or themes — colloquially referred to as subreddits. Previous research has explored the relationship between news coverage and Reddit engagement. (Gozzi et al. 2020) demonstrated that COVID-19 news coverage drove users to comment on Reddit and search for information on Wikipedia, though this effect decreased over time — probably due to media saturation. Other research on Reddit as a site of information has found that fact-checked information lasts longer when a post is deemed true (Bond and Garrett 2023), reinforcing that users rely on Reddit’s discussions when interpreting news. However, other research has scrutinized the credibility of information posted to Reddit — without editorial oversight, users may share biased viewpoints and content from known misinformation sources (Chipidza et al. 2022). Tangled into these findings on social media and news consumption is the role Wikipedia might still play in this process. Studies have found that Reddit users regularly rely on Wikipedia hyperlinks to validate information — especially within the “Today I Learned” or “TIL” subreddit (Moyer et al. 2015; Vincent, Johnson, and Hecht 2018). Nonetheless, access to this previous dataset is no longer feasible — since 2023 Reddit’s API is no longer available for free public use. Researchers wishing to access Reddit data must now submit an application to gain access to the “Reddit4Researchers” beta program — a new approach to partnering with data scientists to balance between data accessibility and user protection (Perez 2024).
@@
For the purposes of this dataset, only post, comment, and subreddit IDs, together with anonymous metadata such as timestamps, score, and extracted Wikipedia links, are shared. All IDs are securely hashed using SHA-256. This measure preserves individuals’ privacy and also enables future researchers to collect and analyze additional data on entries of interest with access to the Reddit4Researchers API by matching against our hashes. This data is stored in the posts and comments tables.
@@
This work was approved by the University of Exeter ethical review procedures (ID 8969882). The dataset is based on public data collected via the Reddit4Researchers and Wikipedia APIs. This research was observational, and no personally identifiable information from Reddit users was included. Data is limited to publicly accessible posts and comments containing Wikipedia links, excluding private and NSFW subreddits. The dataset conform to the FAIR principles (Wilkinson et al. 2016) as follows.
@@
Figure 3: Plot showing the daily average Reddit score of the posts and comments that mention Wikipedia (in text or as a link) over 2020-2023.
@@
For page views, we first deal with 0 counts, typically indicating an article doesn’t yet exist or is deleted by removing them from analysis. We then compute the geometric mean of the relative values, since the distribution is extremely long-tailed. We find that for links in Reddit posts, we observe a 45% increase in page views on the day of posting, and a 6% increase in the week after posting, as compared to the week before posting. For links in Reddit comments, we observe a 45% increase in page views on the day of posting, and a 5% increase in the week after posting, as compared to the week before posting. The similarity between posts and comments in Figure 4 would suggest much of this activity is due to some external stimulus, as one would assume the level of attention to posts vs comments, and the subsequent spillovers to Wikipedia, would be different. This indicates that audiences are not passive news receivers, but consult different online platforms to support their comprehension of current events.
@@
The picture regarding edits is less clear. This is partially due to the fact that many of the linked articles receive no edits in the days before / on / after being posted on Reddit, in line with the findings of Vincent, Johnson, and Hecht (2018). Nevertheless, we do see some indications of increased editing activity in the wake of being posted. We first remove cases where the pages receive no edits over the full time period, then, due to the smaller, discrete edit values, consider the absolute change in number of daily edits. We compare the number of daily edits on the day of posting on Reddit and in the week after posting against that in the week before posting. Considering the arithmetic mean, for links in Reddit posts, we observe an increase of 0.462 daily edits on the day of posting, but a decrease of 0.031 daily edits in the week after posting, as compared to the week before posting. For links in Reddit comments, we observe an increase of 0.127 daily edits on the day of posting, but a decrease of 0.022 daily edits in the week after posting, as compared to the week before posting. The relatively small effect size here, plus high levels of zero-inflation, and the presence of extreme outliers warrants more rigorous analysis.
@@
Both Reddit and Wikipedia are used in different languages. Regarding the use of Wikipedia links on the social media platform, activity is very much English language dominated. 95.8% of Reddit posts with Wikipedia links in the dataset are in English and 93.9% of all linked articles are to English Wikipedia, with the next most links being to German, French, and Spanish Wikipedias (Fig. 5). However, there is substantial cross-lingual linking. 50.85% of links to non-English Wikipedias come from English language Reddit posts. It is also notable how frequently non-English language Reddit posts link to the English language Wikipedia (Fig. 6). The most active other languages are less likely to link to English Wikipedia. An interesting dynamic is observed for Esperanto (eo), the artificially constructed international second language. Esperanto Wikipedia exists, but Esperanto users on Reddit almost exclusively use English Wikipedia—presumably their first language. The use of non-English Wikipedia articles in English Reddit posts may highlight knowledge gaps on the English Wikipedia, or be used to compare accounts in different language editions. The extensive use of English Wikipedia in other language Reddit posts strongly highlights (perceived) knowledge gaps in these languages. Unfortunately, it is likely that this linking behavior weakens the already fragile pipeline of contributions coming from social media, lessening the likelihood of edits being made to the Wikipedia language editions in need of them. These results, and future work, can have wide-ranging implications for the multilingual state of Wikipedia and the web.

### v2

@@ Abstract
The World Wide Web is a complex interconnected digital ecosystem, where information and attention flow between platforms and communities throughout the globe. These interactions co-construct how we understand the world, reflecting and shaping public discourse. Unfortunately, researchers often struggle to understand how information circulates and evolves across the web because platform-specific data is often siloed and restricted by linguistic barriers. To address this gap, we present a comprehensive, multilingual dataset capturing all Wikipedia mentions and links shared in posts and comments on Reddit 2020–2023, excluding those from private and NSFW subreddits. Each linked Wikipedia article is enriched with revision history, page view data, article ID, redirects, and Wikidata identifiers. Through a research agreement with Reddit, our dataset ensures user privacy while providing a query and ID mechanism that integrates with the Reddit and Wikipedia APIs. This enables extended analyses for researchers studying how information flows across platforms. For example, Reddit discussions use Wikipedia for deliberation and fact-checking which subsequently influences Wikipedia content, by driving traffic to articles or inspiring edits. By analyzing the relationship between information shared and discussed on these platforms, our dataset provides a foundation for examining the interplay between social media discourse and collaborative knowledge consumption and production.
@@ Introduction
In addition to Wikipedia, audiences consult a range of online services for news including news websites, apps, and social media (St. Aubin and Liedke 2024; Vraga and Tully 2021). People who include, but do not exclusively rely on, social media in their news diets tend to have higher news media knowledge (Schulz, Fletcher, and Nielsen 2024). Among these social media sites is Reddit, a social media platform where community members share, vote, and comment on content and foster community-based engagement on topics or themes—colloquially referred to as subreddits. Previous research has explored the relationship between news coverage and Reddit engagement. Gozzi et al. (2020) demonstrated that COVID-19 news coverage drove users to comment on Reddit and search for information on Wikipedia, though this effect decreased over time—probably due to media saturation. Further research on Reddit has found that fact-checked information lasts longer when a post is deemed true (Bond and Garrett 2023), reinforcing that users rely on Reddit’s discussions when interpreting news. However, other work has scrutinized the credibility of information posted to Reddit—without editorial oversight, users may share biased viewpoints and content from known misinformation sources (Chipidza et al. 2022). Tangled into these findings on social media and news consumption is the role Wikipedia might still play in this process. Studies have found that Reddit users regularly rely on Wikipedia hyperlinks to validate information—especially within the “Today I Learned” subreddit (Moyer et al. 2015; Vincent, Johnson, and Hecht 2018). Nonetheless, access to this previous dataset is no longer feasible—since 2023 Reddit’s API is no longer available for free public use. Researchers wishing to access Reddit data must now submit an application for access to the “Reddit4Researchers” beta program—a new approach to partnering with data scientists to balance between data accessibility and user protection (Perez 2024).
@@
For the purposes of this dataset, only post, comment, and subreddit IDs, together with anonymous metadata such as timestamps, score, and extracted Wikipedia links, are shared. All IDs are securely hashed using SHA-256, and checked for uniqueness. This measure preserves individuals’ privacy and also enables future researchers to collect and analyze additional data on entries of interest with access to the Reddit4Researchers API by matching against our hashes. This data is stored in the posts and comments tables.
@@
This work was approved by the University of Exeter ethical review procedures (ID 8969882). The dataset is based on public data collected via the Reddit4Researchers and Wikipedia APIs. This research was observational, and no personally identifiable information from Reddit users was included. Data is limited to publicly accessible posts and comments containing Wikipedia links, excluding private and NSFW subreddits. The dataset conforms to the FAIR principles (Wilkinson et al. 2016) as follows.
@@
When we analyze the distributions longitudinally, we find that Wikipedia is referenced about 8000 times per day in comments and that number has remained relatively stable over time. Original posts that refer to Wikipedia have decreased slightly over time going from approximately 300 posts per day to around 200 posts per day. While not fully answerable solely with our dataset, we hypothesize that these shifts might be related to changes in platform content moderation, engagement preferences with both platforms (e.g., mobile vs laptop) or other technological affordances (e.g., Reddit’s transition to more image and video-based participation).
@@
Figure 3: Plot showing the daily average Reddit score of the posts and comments that mention Wikipedia (in text or as a link) over 2020–2023.
@@
Information from Wikipedia being posted to Reddit may be an indicator of what kinds of information people are paying most attention to. Topics of news attention often play out on Wikipedia (Gildersleve, Lambiotte, and Yasseri 2023), but these attention patterns—sometimes together with the Wikipedia articles to help contextualize the events—also move throughout the internet. Much like the Gozzi et al. (2020) findings, our data indicates that people trying to make sense of news as it unfolds go between both platforms. This may be reflected in the Wikipedia page view and editing patterns of the posted articles. The posting of articles on Reddit may itself also drive interest and editing activity towards Wikipedia.
@@
For page views, we first deal with 0 counts, typically indicating an article doesn’t yet exist or is deleted by removing them from analysis. We then compute the geometric mean of the relative values, since the distribution is extremely long-tailed. We find that for links in Reddit posts, we observe a 45% increase in page views on the day of posting, and a 6% increase in the week after posting, as compared to the week before posting. For links in Reddit comments, we observe a 45% increase in page views on the day of posting, and a 5% increase in the week after posting, as compared to the week before posting.
The similarity between posts and comments in Figure 4 would suggest much of this activity is due to some external stimulus, as one would assume the level of attention to posts vs comments, and the subsequent spillovers to Wikipedia, would be different. This indicates that audiences are not passive news receivers, but consult different online platforms to support their comprehension of current events.
@@
Edits: The picture regarding edits is less clear. This is partially due to the fact that many of the linked articles receive no edits in the days before / on / after being posted on Reddit, in line with the findings of Vincent, Johnson, and Hecht (2018). Nevertheless, we do see some indications of increased editing activity in the wake of being posted. We first remove cases where the pages receive no edits over the full time period, then, due to the smaller, discrete edit values, consider the absolute change in number of daily edits. We compare the number of daily edits on the day of posting on Reddit and in the week after posting against that in the week before posting. Considering the arithmetic mean, for links in Reddit posts, we observe an increase of 0.462 daily edits on the day of posting, but a decrease of 0.031 daily edits in the week after posting, as compared to the week before posting. For links in Reddit comments, we observe an increase of 0.127 daily edits on the day of posting, but a decrease of 0.022 daily edits in the week after posting, as compared to the week before posting. The relatively small effect size here, plus high levels of zero-inflation, and the presence of extreme outliers warrants more rigorous analysis.
@@
Both Reddit and Wikipedia are used in different languages. Regarding the use of Wikipedia links on the social media platform, activity is very much English language dominated. 95.8% of Reddit posts with Wikipedia links in the dataset are in English and 93.9% of all linked articles are to English Wikipedia, with the next most links being to German, French, and Spanish Wikipedias (Fig. 5). However, there is substantial cross-lingual linking. 50.85% of links to non-English Wikipedias come from English language Reddit posts. It is also notable how frequently non-English language Reddit posts link to the English language Wikipedia (Fig. 6). The most active other languages are less likely to link to English Wikipedia. An interesting dynamic is observed for Esperanto (eo), the artificially constructed international second language. Esperanto Wikipedia exists, but Esperanto users on Reddit almost exclusively use English Wikipedia—presumably their first language.
The use of non-English Wikipedia articles in English Reddit posts may highlight knowledge gaps on the English Wikipedia, or be used to compare accounts in different language editions. The extensive use of English Wikipedia in other language Reddit posts strongly highlights (perceived) knowledge gaps in these languages. Unfortunately, it is likely that this linking behavior weakens the already fragile pipeline of contributions coming from social media, lessening the likelihood of edits being made to the Wikipedia language editions in need of them. These results, and future work, can have wide-ranging implications for the multilingual state of Wikipedia and the web.